OpenStudy (anonymous):

Solve a0 = 1; a1 = 6 and an = 6an-1 - 9an-2 + n; for n greater than or equal to 2? Find Recurrence Formula....

OpenStudy (anonymous):

You have the recurrence formula. You need a[n] in terms of n. Right

OpenStudy (anonymous):

This is is a difference equation problem

OpenStudy (anonymous):

The homogeneous equation is \[ a_n = 6 a_{n-1} - 9 a_{n-2} \]

OpenStudy (anonymous):

The characteristic equation is \[p(t)=t^2-6 t+9=(t-3)^2 \] It has a double root r=3. The solution of the homogeneous equation is \[ a[n_]= ( A + B\ n) 3^n \]

OpenStudy (anonymous):

I mean \[ a_n= ( A + B \ n )3^n \]

OpenStudy (anonymous):

Now you have to find a particular solution of the non homogeneous equation of the form F n + G. You solve and you find \[ F=\frac 1 4 \\ G=\frac 3 4\\ \]

OpenStudy (anonymous):

The general solution is \[ a_n= ( A + B \ n )3^n +\frac n 4 + \frac 3 4 \]

OpenStudy (anonymous):

To find A and B, use the initial conditions.

OpenStudy (anonymous):

You find \[ A= \frac 14 \\ B= \frac {17}{12} \]

OpenStudy (anonymous):

The final answer is \[ a_n=3^n \left(\frac{17 n}{12}+\frac{1}{4}\right)+\frac{n}{4}+\frac{3}{4} \]

OpenStudy (anonymous):

Thank you so so much! This helps more than you could know. The way you described things really makes sense to me too. You're a lifesaver :)